A jet flew from Tokyo to Bangkok, a distance of 4800km. On the return trip, the speed was decreased by 200km/h. If the difference in the times of the flights was 2 hours, what was the jet's speed from Bangkok to Tokyo? (13 marks)
Answer by stanbon(48558)

A jet flew from Tokyo to Bangkok, a distance of 4800km. On the return trip, the speed was decreased by 200km/h. If the difference in the times of the flights was 2 hours, what was the jet's speed from Bangkok to Tokyo? (13 marks)
----------------------------------------------
T to B DATA:
distance = 4800 km ; rate = x km/h ; time = d/r = 4800/x hrs
----------------------------------------------
B to T DATA
distance = 4800 km ; rate = x-200 km/h ; time = 4800/(x-200) hrs
----------------------------------------------
Equation:
slower time - faster time = 2 hrs
4800/(x-200) 4800/x = 2
4800x - 4800(x-200) = 2x(x-200)
960000 = 2x^2-400x
x^2 - 200x - 480000 = 0
(x - 800)(x + 600) = 0
x = 800 km/h (speed from B to T)
======================================
Cheers,
Stan H.